OpenStudy (anonymous):

solve the equations: coshx - 3sinhy = 0 2sinhx + 6coshy = 5

OpenStudy (anonymous):

not sure the best approach ..

OpenStudy (amistre64):

cramers rule

OpenStudy (anonymous):

not been a mention of that in this chapter of the book

OpenStudy (anonymous):

@amistre cramer or write as exponentials? i am not sure

OpenStudy (anonymous):

i tried as exponentials maybe i need to persevere with that

OpenStudy (amistre64):

\[{{ax\ by\ |\ n}\choose{cx\ dy\ |\ m}}\] \[y=\frac{am-cn}{ad-cb}\] \[x=\frac{bm-dn}{-(ad-cb)}\]

OpenStudy (anonymous):

is cramer post A level?

OpenStudy (amistre64):

i got no idea what level cramer is :)

OpenStudy (amistre64):

coshx - 3sinhy = 0 2sinhx + 6coshy = 5 \[\frac{hmmm}{6cosh(y)cosh(x)}\] never tried it with these things tho

OpenStudy (amistre64):

im sure there are identities involved but i aint got good experience with the hyperbolic stuff: i recall that: cos^2 - sin^2 = 1

OpenStudy (anonymous):

i suspect its not the technique here, as i've never hear of cramers mentioned

OpenStudy (anonymous):

i've tried that route it didn't get me anywhere either

OpenStudy (anonymous):

i think i'll go back to exponential method see if it will yield like that

OpenStudy (amistre64):

the exp route is turning them to e^(...) right?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

hmmm, this might be useful for a chk then: http://www.wolframalpha.com/input/?i=coshx+-+3sinhy+%3D+0+and+2sinhx+%2B+6coshy+%3D+5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got that but it didn't help would it help to write \[y=\sinh^{-1}(\frac{\cosh(x)}{3})\]

OpenStudy (anonymous):

might do i'll check it out

OpenStudy (anonymous):

multiplied first one by 2, added to second and got \[2e^x+6e^{-y}=5\] but i am not sure that helps either

OpenStudy (amistre64):

\[cosh(x)=\frac{e^x+e^{-x}}{2};\ sinh(x)=\frac{e^x-e^{-x}}{2}\] \begin{array}l cosh(x) - 3sinh(y) = 0\\ 2sinh(x) + 6cosh(y) = 5\end{array} \begin{array}l \frac{e^x+e^{-x}}{2} - 3\frac{e^y-e^{-y}}{2} = 0\\ 2\frac{e^x-e^{-x}}{2} + 6\frac{e^y+e^{-y}}{2} = 5\end{array} \begin{array}l e^x+e^{-x} - 3e^y+3e^{-y} = 0\\ 2e^x-2e^{-x} + 6e^y+6e^{-y} = 10\end{array} \begin{array}l -2e^x-2e^{-x} +6e^y-6e^{-y} = 0\\ 2e^x-2e^{-x} + 6e^y+6e^{-y} = 10\end{array} \[-4e^{-x} +12e^y=10\] \[-2e^{-x} +6e^y=5\] would it be simpler to try a substitution? \[cosh(x)=3sinh(y)\]

OpenStudy (anonymous):

so maybe \[y=\frac{arcsinh(x)}{3}\]

OpenStudy (anonymous):

then \[2\cos(x)+6\cosh(\frac{arcsinh(x)}{3})=5\] but then i am stuck again

OpenStudy (anonymous):

meant \[2\sinh(x)+6\cosh(\frac{arcsinh(x)}{3})=5\]